OpenStudy (anonymous):

(t^-4)^-9t^2 Giving out medals & will fan you

OpenStudy (anonymous):

@RadEn Can you help?

OpenStudy (anonymous):

Subtract t from both sides: 4 + t - t = 9t - t 4 = 8t Divide both sides by 8: 4/8 = 8t/8 1/2 = t t = 1/2.

OpenStudy (anonymous):

The possible answers are t^26 t^-26 t^38 t^-38

OpenStudy (raden):

can you give the parantheses, please to make sure

OpenStudy (raden):

is it like this : (t^-4)^-9 * t^2 ?

OpenStudy (anonymous):

It is not written that way, but I believe that is implied.

OpenStudy (raden):

yeah, i think it is (t^-4)^-9 * t^2 first, use the property of exponent : (a^m)^n = a^(m * n)

OpenStudy (raden):

so, it can be written as (t^-4)^-9 * t^2 = t^36 * t^2 agree ?

OpenStudy (anonymous):

You would add & I believe you would come up with t^39 is what my calculation told me.

OpenStudy (anonymous):

t^38**

OpenStudy (raden):

yeah, that's right. first times both powers (in-out oxponents) then add the powers. yes the last answer is t^38

OpenStudy (anonymous):

Thank you for your help! I do have one other question if your up for it?

OpenStudy (raden):

ok, next question :)

OpenStudy (anonymous):

Radio signals travel at a rate of 3*10^8 meters per second. How many seconds would it take for a radio signal to travel from a satellite to the surface of Earth if the satellite is orbiting at a height of 4.2*10^7 meters? 1.26*10^16 seconds 1.26*10^15 seconds 1.4*10^1 seconds 1.4*10^-1 seconds

OpenStudy (raden):

time (t) = 4.2 * 10^7 divided by 3*10^8

OpenStudy (raden):

just simplies that

OpenStudy (anonymous):

1.4x10^15! :D

OpenStudy (raden):

4.2/3 = 1.4 (ok, this is right) but for 10^7 : 10^8, use the property of exponent this a^m : a^n = a^(m-n)

OpenStudy (raden):

so, 10^7 : 10^8 = 10^(7-8) = 10^-1

OpenStudy (anonymous):

Okay. hah Thanks alot dude you really were a life saver! I will be fanning you too :)

OpenStudy (raden):

you're welcome ;)